# Gradient-Free Adversarial Purification with Diffusion Models

### v1

@@ Abstract
Adversarial training and adversarial purification are two effective and practical defense methods to enhance a model’s robustness against adversarial attacks. However, adversarial training necessitates additional training, while adversarial purification suffers from low time efficiency. More critically, current defenses are designed under the perturbation-based adversarial threat model, which is ineffective against the recently proposed unrestricted adversarial attacks. In this paper, we propose an effective and efficient adversarial defense method that counters both perturbation-based and unrestricted adversarial attacks. Our defense is inspired by the observation that adversarial attacks are typically located near the decision boundary and are sensitive to pixel changes. To address this, we introduce adversarial anti-aliasing to mitigate adversarial modifications. Additionally, we propose adversarial super-resolution, which leverages prior knowledge from clean datasets to benignly recover images. These approaches do not require additional training and are computationally efficient without calculating gradients. Extensive experiments against both perturbation-based and unrestricted adversarial attacks demonstrate that our defense method outperforms state-of-the-art adversarial purification methods.
1 Introduction
@@
Deep learning models have demonstrated remarkable performance across various tasks [14, 22, 36]. With the rapid advancement and widespread deployment of these models, their security and robustness are garnering increasing attention.
@@
It is widely recognized that deep learning models are highly vulnerable to adversarial attacks [23, 3]. These attacks are performed by adding imperceptible perturbations to clean images. The perturbed images, known as adversarial examples, can deceive trained deep learning classifiers with high confidence while appearing natural and realistic to human observers. To mitigate adversarial attacks and ensure the stability of deep learning models, adversarial training [23, 13] has been developed. This approach aims to defend against adversarial attacks by training the classifier with adversarial examples. However, adversarial training tends to perform poorly against unknown attacks.
Recently, with the development of diffusion models [11, 26], adversarial purification [24, 29] has shown promising defense performance by recovering the adversarial examples to clean images. These works adopt the diffusion model’s reverse generation process to gradually remove the Gaussian noise from the forward process and the adversarial perturbations. Nevertheless, these methods require heavy computational resources during the purification, which may not be practical in real-time scenarios.
Diffusion models also facilitate stronger unrestricted adversarial attacks [5, 9, 6]. These unrestricted adversarial examples (UAEs) are generated through the reverse generation process by incorporating adversarial guidance. Unlike traditional perturbation-based adversarial attacks, UAEs exhibit superior attack performance against current defenses due to their distinct threat models. These attacks pose a new threat to the development of deep learning models and urgently need to be addressed. Even worse, existing defenses have merely covered the discussion against UAEs.
@@
Figure 1: The proposed adversarial defense pipeline. We give an adversarial example of “cock” class with AutoAttack $\ell_{\text{inf}}=8/255$ on ImageNet dataset. Adversarial anti-aliasing aims to eliminate adversarial perturbations, while adversarial super-resolution seeks to restore benign images from blurred adversarial examples using prior knowledge from the clean dataset.
@@
Our defense employs zero-shot adversarial purification by extracting the “semantic shape” information from images without the image details, as illustrated in Figure 1. Specifically, we use adversarial anti-aliasing with specialized filters to blur the detailed adversarial modifications in the adversarial examples. Following this, we apply adversarial super-resolution to the anti-aliased adversarial examples, upscaling the blurred images using details from pre-trained clean super-resolution diffusion models. These two methods are time-efficient and do not require any modifications to the original models. To demonstrate the effectiveness of our proposed defense, we further validate its performance by using the upscaled adversarial examples as input for adversarial purification. Experiments on various datasets show that our defense outperforms state-of-the-art adversarial defenses in adversarial purification.
@@
  • We introduce various zero-shot and gradient-free defense strategies that preserve the semantic information of adversarial examples while eliminating adversarial modifications. These strategies include adversarial anti-aliasing for “semantic” extraction and adversarial super-resolution for incorporating benign priors and recovering benign details from adversarial examples.
@@
2 Background
2.1 Adversarial Training
Adversarial training (AT) is one of the most practical methods for enhancing a model’s robustness against adversarial attacks. It involves training the model with both benign and adversarial data simultaneously during the training phase. However, robustness against unseen attacks remains a significant challenge that affects the defense performance of traditional adversarial training [23]. To address this, Gowal et al. [13] and Rebuffi et al. [25] have incorporated generated and augmented data to improve generalization by increasing data diversity. In addition to leveraging diverse data, refining the objective formulation of AT has also proven effective. By considering model weights, a wide range of adversarial training methods [35, 17] have been proposed.
2.2 Adversarial Purification
Adversarial purification aims to eliminate adversarial perturbations in adversarial examples without requiring the re-training of deep learning models. These methods leverage the generative capabilities of generative models. Previous works utilizing generative adversarial networks (GANs) [28] and score-based matching models [31, 37] have demonstrated state-of-the-art performance compared to adversarial training. With the advent of diffusion models, Nie et al. [24] discovered that diffusion-based adversarial purification methods outperform previous approaches in recovering clean images. However, finding the optimal generation steps for diffusion-based adversarial purification remains challenging. Additionally, adversarial images can negatively impact the reverse generation process of diffusion models. To address these issues, several works [33, 20, 29] have proposed various solutions to enhance the performance of adversarial purification.
3 Preliminary
3.1 Threat Model
@@
$A_{\text{AE}}\triangleq\{x_{\text{adv}}=x+\delta|y\neq f(x),x\in D,|\delta|\leq\epsilon\}$ (1)
where $\delta$ is the adversarial perturbation, $f(\cdot)$ is the target model, $D$ is the clean dataset, and $\epsilon$ is the perturbation norm constraint.
@@
$A_{\text{UAE}}\triangleq\{x_{\text{adv}}\in\mathcal{G}(z_{\text{adv}},y)|y\neq f(x)\}$ (2)
where $\mathcal{G}$ is the generation model, $z_{\text{adv}}$ is the latent code for generation.
@@
Figure 2: The comparisons of state-of-the-art diffusion-based adversarial purification pipelines. We mark the defense process in blue to represent time-consuming approaches. We use red font to indicate non-purified adversarial input.
3.2 Diffusion-Based Adversarial Purification
The diffusion model [16] learns to recover the image from the denoising-like process, i.e., reverse generation process. The reverse generation process takes $T$ time steps to obtain a sequence of noisy data $\{x_{T-1},\dots,x_{1}\}$ and get the data $x_{0}$ at the last step. Specifically, it can be formulated as:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1}:\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ (3)
The forward diffusion process is where we iteratively add Gaussian noise to the data for training the diffusion model to learn $p_{\theta}(x_{t-1}|x_{t})$. It is defined as:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t}:\sqrt[]{\sigma_{t}}x_{t-1},(1-\sigma_{t})\textbf{I})$ (4)
@@ 3.2 Diffusion-Based Adversarial Purification
[…]
=${\sigma_{t^{*}}}(x+\delta)+\sqrt[]{1-\sigma_{t^{*}}}\varepsilon$
where $\varepsilon$ is the Gaussian noise ${\varepsilon}\sim\mathcal{N}(0,\textbf{I})$. After we obtain the optimal $t^{*}$, we can utilize the reverse generation process over $x_{\text{adv}}$ to recover the clean $x$.
@@
Song et al. [29] utilized the whole reverse generation process with $T$ time step; they used adversarial sample $x_{\text{adv}}$ as guidance rather than an intermediate time step state. At each time step $t$, the guidance is added to the $x_{t}$ after the original reverse generation process and can be formulated as:
$\nabla_{{x}}\log p(x_{\text{adv}}|x_{t};t)=-R_{t}\nabla_{{x}_{t}}d(\hat{x}_{t},x_{\text{adv}})$ (6)
where $R_{t}$ is the scale factor at $t$ time step, $d(\cdot)$ is the distance measurement, and $\hat{x}_{t}$ is the estimation for $x_{0}$ at $t$ time step. The $\hat{x}_{t}$ is defined as:
[…]
@@
4 Methodology
4.1 Motivation
Figure 4: The example of proposed adversarial super-resolution. Our method achieves similar adversarial purification without any gradient calculation of diffusion models.
@@ 4.1 Motivation
4.2 Perturbation-Isolated Adversarial Purification
Perturbation-based adversarial examples are precisely calculated based on the gradient of the loss function, whereas unrestricted adversarial examples are sampled near the decision boundary. Despite employing different threat models, both types of attacks produce adversarial examples that are sensitive to pixel changes. Since adversarial examples are designed to be imperceptible compared to clean images, the semantic shapes of objects within the images should correspond to their original labels. Therefore, our defense strategy focuses on extracting the semantic shapes from the adversarial examples and eliminating the adversarial pixel-level details.
4.2.1 Adversarial Anti-Aliasing

### v2

@@ Abstract
Adversarial training and adversarial purification are two widely used defense strategies for enhancing model robustness against adversarial attacks. However, adversarial training requires costly retraining, while adversarial purification often suffers from low efficiency. More critically, existing defenses are primarily designed under the perturbation-based adversarial threat model, which is ineffective against recently introduced unrestricted adversarial attacks. In this paper, we propose an effective and efficient defense framework that counters both perturbation-based and unrestricted adversarial attacks. Our approach is motivated by the observation that adversarial examples typically lie near the decision boundary and are highly sensitive to pixel-level perturbations. To address this, we introduce adversarial anti-aliasing, a preprocessing technique that mitigates adversarial noise by reducing the magnitude of pixel-level perturbations. In addition, we propose adversarial super-resolution, which leverages prior knowledge from clean datasets to benignly restore high-quality images from adversarially degraded ones. Unlike image synthesis methods that generate entirely new images, adversarial super-resolution focuses on image restoration, making it more suitable for purification. Importantly, both techniques require no additional training and are computationally efficient since they do not rely on gradient computations. To further improve robustness across diverse datasets, we introduce a contrastive learning–based adversarial deblurring fine-tuning method. By incorporating adversarial priors during fine-tuning on the target dataset, this method enhances purification effectiveness without the need to retrain diffusion models. Extensive experiments demonstrate that our defense consistently outperforms state-of-the-art adversarial purification methods against both perturbation-based and unrestricted adversarial attacks, achieving superior standard accuracy, robust accuracy, and time efficiency. Moreover, our defense framework is highly flexible and can be seamlessly integrated with existing defense strategies for even stronger purification.
Index Terms:
Adversarial defense, adversarial purification, unrestricted adversarial example, diffusion model, anti-aliasing, super resolution.
I Introduction
@@
Deep learning models have demonstrated remarkable performance across various tasks [1, 2, 3]. With the rapid advancement and widespread deployment of these models, their security and robustness are garnering increasing attention from the research community.
@@
It is widely recognized that deep learning models are highly vulnerable to adversarial attacks [4, 5]. These attacks are performed by adding imperceptible perturbations to clean images. The perturbed images, known as adversarial examples, can deceive trained deep learning classifiers with high confidence while appearing natural and realistic to human observers. To mitigate adversarial attacks and ensure the stability of deep learning models, adversarial training [4, 6] has been developed. This approach defends against adversarial attacks by training the classifier with adversarial examples. However, adversarial training suffers from two major limitations. First, it incurs substantial computational overhead. Unlike standard training, which requires only a single backward propagation per step, adversarial training necessitates generating adversarial examples through multiple rounds of gradient backpropagation. As a result, the overall training cost is several times higher than that of standard training. Second, adversarial training exhibits poor transferability to unseen attacks. It is typically effective only against the specific attacks included during training. To mitigate newly emerging adversarial threats, the model must undergo additional retraining, further compounding its inefficiency.
Adversarial purification [7, 8] seeks to remove adversarial perturbations from adversarial examples and restore clean images by introducing a pre-processing model before the target model. Unlike adversarial training, purification does not require retraining the target model and thus significantly reduces training overhead. Early approaches employed denoising networks or generative models, such as generative adversarial networks (GANs). However, these methods often suffered from overfitting or the limited generative capacity of the models. With the recent advances in diffusion models [9, 10], diffusion-based adversarial purification [11, 12, 13] has demonstrated strong defense performance by effectively recovering adversarial examples to clean images. These methods leverage the reverse generation process of diffusion models, which gradually removes both Gaussian noise introduced in the forward process and adversarial perturbations. Diffusion-based purification is particularly suitable for this task, as the denoising mechanism inherent in diffusion training and sampling can neutralize adversarial noise through injected Gaussian perturbations. Despite their effectiveness, diffusion-based purification methods face a major limitation: the sampling process is computationally expensive and time-consuming, making them impractical for real-time defense scenarios.
Diffusion models also facilitate stronger unrestricted adversarial attacks [14, 15, 16]. These unrestricted adversarial examples (UAEs) are generated through a reverse-generation process guided by adversarial objectives. Unlike traditional perturbation-based attacks, UAEs demonstrate superior effectiveness against current defenses due to their fundamentally different threat model. They present a significant challenge to the development of deep learning models, as they are directly generated by diffusion models rather than being crafted through small perturbations on benign data. As out-of-distribution samples, UAEs can effectively bypass neural networks trained on standard data distributions. Alarmingly, existing defenses provide only limited coverage against these attacks, highlighting the urgent need for new defense strategies specifically targeting UAEs.
@@
Figure 1: The proposed adversarial defense pipeline. We give an adversarial example of “cock” class with AutoAttack $\ell_{\infty}=8/255$ on ImageNet dataset. Adversarial anti-aliasing aims to eliminate adversarial perturbations, while adversarial super-resolution seeks to restore benign images from blurred adversarial examples using prior knowledge from the clean dataset.
@@
Our defense adopts a zero-shot adversarial purification strategy by extracting the semantic shape of images while discarding fine-grained details with adversarial perturbations, as illustrated in Figure 1. Specifically, we employ adversarial anti-aliasing with specialized filters to suppress adversarial perturbations, followed by adversarial super-resolution, which reconstructs high-quality details from pre-trained clean super-resolution diffusion models. Both methods are time-efficient and require no modifications to the original models. To further enhance performance across diverse target datasets, we introduce a contrastive learning-based adversarial deblurring fine-tuning method. This approach leverages our adversarial anti-aliasing to construct paired low-quality and high-quality images, enabling the model to learn dataset-specific priors and thereby improve purification effectiveness on various trained target models. Additionally, we incorporate adversarial contrastive learning into the deblurring task, equipping the diffusion model with adversarial-aware priors to strengthen defense robustness. We validate the effectiveness of our defense by applying the purified adversarial examples as inputs to existing adversarial purification methods, demonstrating substantial improvements in their defense performance. Experiments across multiple datasets confirm that our approach not only outperforms state-of-the-art adversarial defenses in purification quality but also achieves an average $\times 44$ speed-up in time efficiency compared to existing diffusion-based defenses.
@@
  • We introduce various zero-shot and gradient-free defense strategies that preserve the semantic information of adversarial examples while eliminating adversarial modifications with theoretical supports. These strategies include adversarial anti-aliasing for “semantic” extraction and adversarial super-resolution for incorporating benign priors and recovering benign details from adversarial examples.
  • We present a contrastive learning-based adversarial deblurring fine-tuning method, which learns dataset-specific and adversarial-aware priors to further enhance the performance of our adversarial purification across diverse target datasets.
@@
II Background
II-A Adversarial Training
Adversarial training (AT) is among the most practical and widely studied methods for enhancing model robustness against adversarial attacks [17]. The central idea is to expose the model to both benign and adversarial examples during training, enabling it to learn representations that are resilient to malicious perturbations. Madry et al. [4] provided a comprehensive formulation of adversarial training as a robust optimization problem, establishing a strong theoretical foundation. Building on this, Kannan et al. [18] improved the performance of AT by encouraging similarity in the logits of paired examples, thereby stabilizing learning. Despite these advances, a key limitation of adversarial training lies in its trade-off between robustness and standard accuracy: while it enhances robustness, the introduction of adversarial examples during training often reduces clean accuracy. Zhang et al. [19] addressed this issue by designing specialized loss functions that explicitly balance clean and robust accuracy, leading to significant improvements in both. Nevertheless, robustness against unseen or out-of-distribution attacks remains a major challenge for traditional AT. To overcome this limitation, Song et al. [20] incorporated domain adaptation techniques to enhance the generalization ability of adversarially trained models, while Gowal et al. [6] showed that leveraging generated or even unrealistic synthetic data can further strengthen robustness. Another well-documented issue in adversarial training is robust overfitting, where models overfit to the specific adversarial examples seen during training, thus weakening generalization to novel attacks. Rebuffi et al. [21] addressed this by incorporating augmented data, which increased diversity and mitigated overfitting. In addition to enriching training data, refining the training objective has proven effective in improving generalization. Wu et al. [22] proposed Adversarial Weight Perturbation (AWP), which extends perturbations not only to the inputs but also to the network weights, thereby improving robustness against a broader range of attacks. More recently, Jin et al. [23] introduced stochasticity into adversarial training by applying a robustness loss function over randomized weights, achieving gains in both clean and robust accuracy.
II-B Adversarial Purification
Adversarial purification aims to remove adversarial perturbations from adversarial examples without requiring the retraining of deep learning models. Early approaches [24, 25] adopted image transformation techniques like JPEG compression to suppress adversarial noise, but these methods were effective only against small perturbations and lacked robustness against stronger attacks. To improve purification strength, subsequent works leveraged the generative capabilities of deep generative models. Samangouei et al. [7] employed Generative Adversarial Networks (GANs) to sample clean images close to adversarial counterparts by training a defense GAN. Score-based generative models [26] also demonstrated strong purification capabilities. For example, Yoon et al. [27] introduced an Energy-Based Model combined with Markov Chain Monte Carlo (MCMC) sampling for purification, further improving sampling efficiency through denoising score matching. With the emergence of diffusion models [28, 9], a new paradigm of adversarial purification was introduced. Nie et al. [11] showed that diffusion-based purification significantly outperforms prior approaches by exploiting the reverse generation process. Their method adopting the diffusion reverse sampling process with optimal diffusion steps to remove the adversarial perturbations within the sampled Gaussian noise. However, finding the optimal generation steps for diffusion-based adversarial purification remains challenging. To address this limitation, Lee et al. [29] proposed a gradual noise-scheduling strategy with tailored forward-step schedules, improving purification stability. Bai et al. [30] introduced contrastive guidance to refine the sampling process and enhance purification performance. Lin et al. [31] incorporated adversarial training into diffusion models, further strengthening their defensive capability. Another challenge lies in the vulnerability of diffusion’s reverse generation process to adversarial images. Wang et al. [32] tackled this problem using a guided diffusion approach, which eliminates the need for optimal step selection by initiating purification from random Gaussian noise rather than adversarial inputs. Building on this, Song et al. [12] introduced proposed two guidance, long-range and short-range guidance, to enhance the adversarial purification with the guided diffusion model.
Despite these advances, existing purification methods face two major limitations: (1) low time efficiency, particularly for gradient-based guidance methods, and (2) poor effectiveness against unrestricted adversarial examples (UAEs). These challenges highlight the need for more efficient and generalizable adversarial purification techniques.
Figure 2: The comparisons of state-of-the-art diffusion-based adversarial purification pipelines. We mark the defense process in blue to represent time-consuming approaches. We use red font to indicate non-purified adversarial examples with adversarial perturbations.
III Preliminary
III-A Threat Model
@@
$A_{\text{AE}}\triangleq\{\bm{x}_{\text{adv}}=\bm{x}+\bm{\delta}|y\neq f(\bm{x}),\bm{x}\in\mathcal{D},|\bm{\delta}|\leq\kappa\}$ (1)
where $\bm{\delta}$ is the adversarial perturbation, $f(\cdot)$ is the target model, $\mathcal{D}$ is the clean dataset, and $\kappa$ is the perturbation norm constraint.
@@
$A_{\text{UAE}}\triangleq\{\bm{x}_{\text{adv}}\in\mathcal{G}(\bm{z}_{\text{adv}},y)|y\neq f(\bm{x})\}$ (2)
where $\mathcal{G}$ is the generation model, $\bm{z}_{\text{adv}}$ is the latent code for generation.
@@
III-B Diffusion-Based Adversarial Purification
The diffusion model [28] learns to recover the image from the denoising-like process, i.e., reverse generation process. The reverse generation process takes $T$ time steps to obtain a sequence of noisy data $\{\bm{x}_{T-1},\dots,\bm{x}_{1}\}$ and get the data $\bm{x}_{0}$ at the last step. Specifically, it can be formulated as:
$p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1}:\mu_{\theta}(\bm{x}_{t},t),\Sigma_{\theta}(\bm{x}_{t},t))$ (3)
The forward diffusion process is where we iteratively add Gaussian noise to the data for training the diffusion model to learn $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$. It is defined as:
$q(\bm{x}_{t}|\bm{x}_{t-1})=\mathcal{N}(\bm{x}_{t}:\sqrt[]{\sigma_{t}}\bm{x}_{t-1},(1-\sigma_{t})\textbf{I})$ (4)
@@ III-B Diffusion-Based Adversarial Purification
$\bm{x}_{t^{*}}$ $=\sqrt[]{\sigma_{t^{*}}}\bm{x}_{\text{adv}}+\sqrt[]{1-\sigma_{t^{*}}}\bm{\varepsilon}$ (5)
$=\sqrt[]{\sigma_{t^{*}}}(\bm{x}+\bm{\delta})+\sqrt[]{1-\sigma_{t^{*}}}\bm{\varepsilon}$
where $\bm{\varepsilon}$ is the Gaussian noise $\bm{\varepsilon}\sim\mathcal{N}(0,\textbf{I})$. After we obtain the optimal $t^{*}$, we can utilize the reverse generation process over $\bm{x}_{\text{adv}}$ to recover the clean $\bm{x}$.
@@
Song et al. [12] utilized the whole reverse generation process with $T$ time step; they used adversarial sample $\bm{x}_{\text{adv}}$ as guidance rather than an intermediate time step state. At each time step $t$, the guidance is added to the $\bm{x}_{t}$ after the original reverse generation process and can be formulated as:
$\nabla_{\bm{x}}\log p(\bm{x}_{\text{adv}}|\bm{x}_{t};t)=-R_{t}\nabla_{\bm{x}_{t}}d(\bm{\hat{x}}_{t},\bm{x}_{\text{adv}})$ (6)
where $R_{t}$ is the scale factor at $t$ time step, $d(\cdot)$ is the distance measurement, and $\bm{\hat{x}}_{t}$ is the estimation for $\bm{x}_{0}$ at $t$ time step. The $\bm{\hat{x}}_{t}$ is defined as:
$\bm{\hat{x}}_{t}=\frac{\bm{x}_{t}-\sqrt[]{1-\sigma_{t}}s_{\theta}(\bm{x}_{t})}{\sqrt[]{\sigma_{t}}}$ (7)
@@
IV Gradient-Free Adversarial Super-Resolution Purification
IV-A Motivation
@@
Figure 4: The proposed adversarial purification pipeline. Our proposed adversarial purification framework consists of two defense stages. In the first stage, we apply adversarial anti-aliasing as a pre-processing technique to reduce the magnitude of adversarial perturbations while preserving the semantic structure of the image through controlled pixel degradation. Specifically, we perform channel-wise 1D convolution on each RGB channel of the adversarial input, using in-place operations to preserve the original image resolution. The processed RGB channels are then combined to form the anti-aliased image. In the second stage, we apply adversarial super-resolution using a pretrained diffusion model. The purification is performed via the reverse generation process of the super-resolution (SR) diffusion model, which enhances image quality while mitigating residual adversarial noise. To further improve robustness, we introduce deblurring fine-tuning of the SR model, as detailed in Section V. The final purified image is obtained as the output of this SR diffusion process.
IV-B Perturbation-Isolated Adversarial Purification
Perturbation-based adversarial examples are precisely calculated based on the gradient of the loss function, whereas unrestricted adversarial examples are sampled near the decision boundary. Despite employing different threat models, both types of attacks produce adversarial examples that are sensitive to pixel changes. Since adversarial examples are designed to be imperceptible compared to clean images, the semantic shapes of objects within the images should correspond to their original labels. Therefore, our defense strategy focuses on extracting the semantic shapes from the adversarial examples and eliminating the adversarial pixel-level details, as illustrated in Figure 4.
IV-B1 Adversarial Anti-Aliasing